

********************************************************************
UARS WEEKLY STATUS REPORT - 30 JUL 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 July 2004 (Orbit 70308, GMT Day 199) through 23
July 2004 (Orbit 70412, GMT Day 205).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 203/1230 70375 DAY ONLY TO FULLTIME
SUSIM 203/1552 70378 POWER ON
PEM 204/1407 70391 OFF TO DAY ONLY
A Forward-to-Reverse Yaw Around maneuver (number 131) was successfully
executed in Orbit 70317 beginning 199/ 15:04:25Z and completing at
199/ 15:41:35Z on 17 July 2004. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 57.89 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.08 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.51 and 20.04 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 18
August 2004 (DOY 231).
The beta angle ranged between the angles of 01.3 degrees to 19.2
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 39 degrees on 01 August 2004 (DOY 214).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.96 degrees C. Additional
battery performance data for each battery on 16 July and 23 July 2004
are provided in Appendix D.
The clock error ranged from 10.8 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
SOLSTICE = 01
SUSIM = 01
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Three (3) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued testing the new version of the Eclipse ground system
software. There are several DR fixes incorporated in the build, so
testing will take some time. The goal for the testing to be complete
is the end of the July. If all testing is successful the new build
will be released to operations.
The new CMS code to allow for SMA events continued to undergo testing
this week. Problems receiving products from FDF, transferring the
schedule from UPS to CMS, and system drive issues hindered progress.
While the system drive issues have been resolved, the other two issues
are still under investigation.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 August 2004 (DOY 231).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 July 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70317 199/1449 04-210 MACS YAW MANEUVER COMMANDS
70326 200/0623 04-211 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70347 201/1457 04-210 MPS VT-6 TO VT-5
70348 201/1638 04-212 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70375 203/1230 04-213 SOLSTICE DAY ONLY TO FULLTIME
70378 203/1552 04-214 SUSIM POWER ON (PARTIAL)
70390 204/1227 04-214 SUSIM POWER ON (COMPLETE)
70391 204/1407 04-215 PEM OFF TO DAY ONLY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 July 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 July 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70326 200/0630 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70348 201/1639 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 July 2004
16 July 2004 - GMT Day 198 Beta = 04.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -05.6/-22.4 +4.51/+3.87 +16.8/-12.0 26.7 1.11 18.8
3 +16.8/-22.4 +3.55/+2.92 +16.0/-11.6 26.7 1.11 18.4
23 July 2004 - GMT Day 205 Beta = 04.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-44.8 +4.19/+3.23 +27.2/-12.8 26.2 1.01 18.5
3 +11.2/-72.8 +2.92/+2.29 +25.6/-12.8 26.2 1.00 18.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
70347 200/1457 VT6 -> VT5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 July 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT 70335 200/2037 30832 Generic Late Acquisition
04-039 FOT 70335 202/0800 30832 Generic Late Acquisition
04-047 FOT 70367 202/2252 N/A WS-4 not responding properly
04-048 FOT 70395 204/1920 36444 Data Dropout
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-043 FOT 70203 192/0155 30824 Data Dropout (18 sec)
04-044 SSPP 70234 194/0229 N/A SSPP File Xfer Error on CMS
04-045 FOT 70258 195/1750 36367 Data Dropout @ White Sands
04-046 FOT 70290 197/2031 36385 Data Dropout Bad Weather @ WS
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT See List July 04 30832 Generic Late Acquisition
04-047 FOT 70367 202/2252 N/A WS-4 not responding
properly
04-048 FOT 70395 204/1920 36444 Data Dropout